### v11

# Degeneration of families of projective hypersurfaces and Hodge conjecture

Abstract
We prove by induction on dimension the Hodge conjecture for smooth complex projective varieties. Let $X$ be a smooth complex projective variety. Then $X$ is birational to a possibly singular projective hypersurface, hence to a smooth projective variety $E_{0}$ which is a component of a normal crossing divisor $E=\cup_{i=0}^{r}E_{i}\subset Y$ which is the singular fiber of a pencil $f:Y\to\mathbb{A}^{1}$ of smooth projective hypersurfaces. Using the smooth hypersurface case ([5] theorem 1), the nearby cycle functor on mixed Hodge module with rational de Rham factor and the induction hypothesis, we prove that a Hodge class of $E_{0}$ is absolute Hodge, more precisely the locus of Hodge classes inside the algebraic vector bundle given the De Rham cohomology of the rational deformation of $E_{0}$ is defined over $\mathbb{Q}$. By [4] theorem 4, we get the Hodge conjecture for $E_{0}$. By the induction hypothesis we also have the Hodge conjecture for $X$.
1 Introduction
The main result of this article is an inductive proof of the Hodge conjecture using the results of [4] and [5]. More precisely, assuming the Hodge conjecture for smooth complex projective varieties of dimension less or equal to $d-1$, the property of satisfying the Hodge conjecture for connected smooth complex projective varieties of dimension $d$ is a birational invariant (proposition 1). Let $X$ be a connected smooth complex projective variety of dimension $d$. then $X$ is birational to a possibly singular hypersurface $Y_{0_{X}}\subset\mathbb{P}^{d+1}_{\mathbb{C}}$ of degree $l_{X}:=deg(\mathbb{C}(X)/\mathbb{C}(x_{1},\cdots,x_{d}))$. Consider
$f:\mathcal{Y}^{\prime}_{\mathbb{C}}=V(\tilde{f})\subset\mathbb{P}^{d+1}_{\mathbb{C}}\times\mathbb{P}(l_{X})\to\mathbb{P}(l_{X})$
the universal family of projective hypersurfaces of degree $l_{X}$, and $0_{X}\in\mathbb{P}(l_{X})$ the point corresponding to $Y_{0_{X}}$. Denote $\Delta\subset\mathbb{P}(l_{X})$ the discriminant locus parametrizing the hypersurfaces which are singular. Let $\bar{S}:=\bar{0_{X}}^{\mathbb{Q}}\subset\mathbb{P}(l_{X})$ the $\bar{\mathbb{Q}}$-Zariski closure of $0_{X}$ inside $\mathbb{P}(l_{X})$. If $Y_{0_{X}}$ is smooth, then $X$ satisfy the hodge conjecture since $Y_{0_{X}}$ satisfy the Hodge conjecture by [5] and $X$ is birational to $Y_{0_{X}}$, using the induction hypothesis. Hence, we may assume that $Y_{0_{X}}$ is singular, that is $0_{X}\in\Delta$. Then, $\bar{S}\subset\Delta$ since $\Delta\subset\mathbb{P}(l_{X})$ is defined over $\mathbb{Q}$. Let $\bar{T}\subset\mathbb{P}(l_{X})$ be an irreducible closed subvariety of dimension $\mathop{\rm dim}\nolimits(\bar{S})+1$ defined over $\bar{\mathbb{Q}}$ such that
$\bar{S}\subset\bar{T}\;\mbox{and}\;T:=\bar{T}\cap(\mathbb{P}(l_{X})\backslash\Delta)\neq\emptyset.$
We take for $\bar{T}=V(f_{1},\cdots,f_{s})\subset\mathbb{P}(l_{X})$ an irreducible complete intersection defined over $\bar{\mathbb{Q}}$ which intersect $\Delta$ properly and such that
  • $V(f_{1})\supset\mathop{\rm sing}\nolimits(\Delta)\cup\bar{S}$
  • for each $r\in[1,\cdots,s]$, $V(f_{1},\cdots,f_{r})\supset\mathop{\rm sing}\nolimits(V(f_{1},\cdots,f_{r-1})\cap\Delta)\cup\bar{S}$ contain inductively the singular locus of the intersection with $\Delta$ and $\bar{S}$.
Then by the weak Lefchetz hyperplane theorem for homotopy groups, $T\subset\mathbb{P}(l_{X})\backslash\Delta$ is not contained in a weakly special subvariety since
$i_{T*}:\pi_{1}(T_{\mathbb{C}})\to\pi_{1}(\mathbb{P}(l_{X})\backslash\Delta),\;\;(\Delta=\Delta_{\mathbb{C}})$
is an isomorphism if $\mathop{\rm dim}\nolimits\bar{T}\geq 2$ (i.e. $\mathop{\rm dim}\nolimits\bar{S}\geq 1)$ and surjective if $\mathop{\rm dim}\nolimits\bar{T}=1$ (i.e. $\mathop{\rm dim}\nolimits\bar{S}=0$). We have then $\bar{T}\cap\Delta=\bar{S}\cup\bar{S}^{\prime}$, where $\bar{S}^{\prime}\subset\bar{T}$ is a divisor. We then consider the family of algebraic varieties defined over $\bar{\mathbb{Q}}$
$\tilde{f}_{X}:\mathcal{Y}_{\bar{T}^{o}}\xrightarrow{\epsilon}\mathcal{Y}^{\prime}_{\bar{T}^{o}}\xrightarrow{f\times_{\mathbb{P}(l_{X})}\bar{T}^{o}}\bar{T}^{o}$
where
  • $\tilde{f}_{X}:\mathcal{Y}\xrightarrow{\epsilon}\mathcal{Y}^{\prime}_{\bar{T}}\xrightarrow{f\times_{\mathbb{P}(l_{X})}\bar{T}}\bar{T}$ with $\epsilon:(\mathcal{Y},\mathcal{E}^{\prime})\to(\mathcal{Y}^{\prime}_{\bar{T}},\mathcal{Y}^{\prime}_{\bar{S}}\cup\mathcal{Y}^{\prime}_{\bar{S}^{\prime}})$ is a desingularization over $\mathbb{Q}$, in particular $\mathcal{E}^{\prime}=\mathcal{E}\cup\mathcal{E}^{\prime\prime}$ with $\mathcal{E}=\cup_{i=0}^{r}\mathcal{E}_{i}:=\tilde{f}_{X}^{-1}(\bar{S})\subset\mathcal{Y}$ which is a normal crossing divisor, in particular $\mathcal{Y}_{T}=\mathcal{Y}^{\prime}_{T}$,
  • $S\subset\bar{S}$ is the open subset such that $S$ is smooth and $\tilde{f}_{X}\times_{\bar{T}}S:\mathcal{E}_{I}:=\cap_{i\in I}\mathcal{E}_{i}\to S$ is smooth projective for each $I\subset[0,\ldots,r]$, $\bar{T}^{o}\subset\bar{T}$ is an open subset such that $\bar{T}^{o}\cap\Delta=S$ and $T^{o}:=\bar{T}^{o}\cap T$ is smooth, for simplicity we denote again $\mathcal{E}$ for the open subset $\mathcal{E}_{S}:=\tilde{f}_{X}^{-1}(S)\subset\mathcal{E}$,
  • $t\in S(\mathbb{C})$ is the point such that $\mathcal{Y}^{\prime}_{\mathbb{C},t}=Y_{0_{X}}$ and $\epsilon_{t}:E_{0}:=\mathcal{E}_{0,\mathbb{C},t}\to Y_{0_{X}}$ is surjective (there is at least one component of $E$ dominant over $Y_{O_{X}}$ since $E$ is dominant over $Y_{O_{X}}$ as $\mathcal{E}$ is dominant over $\mathcal{Y}^{\prime}_{\bar{S}}$), in particular $E_{0}$ is birational to $Y_{0_{X}}$, hence $E_{0}$ is birational to $X$. We denote $E_{i}:=\mathcal{E}_{i,\mathbb{C},t}$ for each $0\leq i\leq r$ and $E:=\mathcal{E}_{\mathbb{C},t}=\cup_{i=0}^{r}E_{i}$.
Using [4] and [5], assuming the Hodge conjecture for smooth projective varieties of dimension less or equal to $d-1$, we then prove the Hodge conjecture for the $E_{i}$, more specifically for $E_{0}$ as follows : Let $p\in\mathbb{Z}$. By the hard Lefchetz theorem on $H^{*}(E_{0})$, we may assume that $2p\leq d$. We then have then a morphism of mixed Hodge module over $S$ (in fact of geometric variation of mixed Hodge structures over $S$) given by the specialization map
$Sp_{\mathcal{E}/\tilde{f}_{X}}:E^{2p}_{Hdg}(\mathcal{E}/S)\to H^{2p}\psi_{S}E_{Hdg}(\mathcal{Y}_{T^{o}}/T^{o})=\psi_{S}E^{2p}_{Hdg}(\mathcal{Y}_{T^{o}}/T^{o}),$
whose de Rham part are morphism of filtered vector bundle over $S$ with Gauss-Manin connexion defined over $\bar{\mathbb{Q}}$. Moreover, we have, see lemma 1, an isomorphism of variation of mixed Hodge structures over $S$
$c_{\mathcal{E}}:=H^{2p-1}(\tilde{f}_{X}\circ i_{\mathcal{E}})_{*}c(\mathbb{Q}^{Hdg}_{\mathcal{E}})\circ c_{V}:E_{Hdg,\mathcal{E}}^{2p-1}(\mathcal{Y}_{\bar{T}^{o}}/\bar{T}^{o})\xrightarrow{\sim}$
$\phi_{S}E_{Hdg,\mathcal{E}}^{2p-1}(\mathcal{Y}_{T^{o}}/T^{o})\xrightarrow{\sim}\ker(Sp_{\mathcal{E}/\tilde{f}_{X}}:E^{2p}_{Hdg}(\mathcal{E}/S)\to\psi_{S}E^{2p}_{Hdg}(\mathcal{Y}_{T^{o}}/T^{o})).$
Consider the exact sequence of variation of mixed Hodge structure over $S$
$0\to W_{2p-1}E_{Hdg}^{2p}(\mathcal{E}/S)\to E_{Hdg}^{2p}(\mathcal{E}/S)\xrightarrow{(i_{i}^{*})_{0\leq i\leq r}:=q}\oplus_{i=0}^{r}E_{Hdg}^{2p}(\mathcal{E}_{i}/S)\to 0$
It induces the exact sequence of presheaves on $S_{\mathbb{C}}^{an}$,
[…]
where $HL^{2p,p}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})\subset E_{DR}^{2p}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})$, $HL^{2p,p}(\mathcal{E}_{i,\mathbb{C}}/S_{\mathbb{C}})\subset E_{DR}^{2p}(\mathcal{E}_{i,\mathbb{C}}/S_{\mathbb{C}})$ are the locus of Hodge classes. Let $\lambda\in F^{p}H^{2p}(E_{0}^{an},\mathbb{Q})$, where we recall $E_{0}=\mathcal{E}_{0,\mathbb{C},t}$ and $E=\mathcal{E}_{\mathbb{C},t}=\cup_{i=0}^{r}E_{i}$. Consider then $\tilde{\lambda}\in H^{2p}(E^{an},\mathbb{Q})$, such that $q(\tilde{\lambda})=\lambda$, and
$Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda})\in Sp_{E/f_{X}}(H^{2p}(E^{an},\mathbb{Q}))\subset i_{t}^{*}\psi_{S}R^{2p}\tilde{f}_{X*}\mathbb{Q}_{\mathcal{Y}_{T^{o},\mathbb{C}}^{an}}.$
Since the monodromy of $R^{2p}\tilde{f}_{X*}\mathbb{Q}_{\mathcal{Y}_{T^{o},\mathbb{C}}}^{an}$ is irreducible as $i_{T^{o}*}:\pi_{1}(T_{\mathbb{C}}^{o})\to\pi_{1}(T_{\mathbb{C}})\to\pi_{1}(\mathbb{P}(l_{X})\backslash\Delta)$ is surjective, we have
$Sp_{\mathcal{E}/\tilde{f}_{X}}(\ker q):=Sp_{\mathcal{E}/\tilde{f}_{X}}(W_{2p-1}E_{DR}^{2p}(\mathcal{E}/S))=0,$
using the fact that there exists a neighborhood $V_{\bar{S}\cup\bar{S}^{\prime}}\subset\bar{T}_{\mathbb{C}}$ of $\bar{S}_{\mathbb{C}}\cup\bar{S}^{\prime}_{\mathbb{C}}$ in $\bar{T}_{\mathbb{C}}$ for the usual complex topology such that the inclusion $i_{\bar{S}_{\mathbb{C}}\cup\bar{S}^{\prime}_{\mathbb{C}}}:\bar{S}_{\mathbb{C}}\cup\bar{S}^{\prime}_{\mathbb{C}}\hookrightarrow V_{\bar{S}\cup\bar{S}^{\prime}}$ admits a retraction $r:V_{\bar{S}\cup\bar{S}^{\prime}}\to\bar{S}_{\mathbb{C}}\cup\bar{S}^{\prime}_{\mathbb{C}}$ which is an homotopy equivalence. The locus of Hodge classes passing through $\lambda$
$\mathbb{V}_{S}^{p}(\lambda):=\mathbb{V}_{S}(\lambda)\cap F^{p}E^{2p}_{DR}(\mathcal{E}_{0,\mathbb{C}}/S_{\mathbb{C}})\subset E^{2p}_{DR}(\mathcal{E}_{0,\mathbb{C}}/S_{\mathbb{C}}),$
inside the De Rham vector bundle of $\tilde{f}_{X}\circ i_{E_{0}}:\mathcal{E}_{0}\to S$ thus satisfy
$\mathbb{V}_{S}^{p}(\lambda)=q(Sp_{\mathcal{E}/\tilde{f}_{X}}^{-1}(\mathbb{V}_{S}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda})))),$ (1)
where
  • $\mathbb{V}_{S}(\lambda)\subset E^{2p}_{DR}(\mathcal{E}_{0,\mathbb{C}}/S_{\mathbb{C}})$, $\mathbb{V}_{S}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))\subset E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T_{\mathbb{C}}^{o})$, are the flat leaves, e.g. $\mathbb{V}_{S}(\lambda):=\pi_{S}(\lambda\times\tilde{S}^{an}_{\mathbb{C}})$ where $\pi_{S}:H^{2p}(E^{an},\mathbb{C})\times\tilde{S}_{\mathbb{C}}^{an}\to E^{2p,an}_{DR}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})$ is the morphism induced by the universal covering $\pi_{S}:\tilde{S}_{\mathbb{C}}^{an}\to S_{\mathbb{C}}^{an}$,
  • $q:=q\otimes\mathbb{C}:E^{2p}_{DR}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})\to E^{2p}_{DR}(\mathcal{E}_{0,\mathbb{C}}/S_{\mathbb{C}})$ is the quotient map.
  • $Sp_{\mathcal{E}/\tilde{f}_{X}}:=Sp_{\mathcal{E}/\tilde{f}_{X}}\otimes\mathbb{C}:E^{2p}_{DR}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})\to i_{S}^{*mod}E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T_{\mathbb{C}}^{o})$.
Now,
  • If $Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda})=0$, we have by lemma 1 applied to $\tilde{f}_{X}:\mathcal{Y}_{\bar{T}^{o}}\to\bar{T}^{o}$, $\tilde{\lambda}=c_{\mathcal{E}}(\tilde{\lambda})$ with $\tilde{\lambda}\in F^{p}H^{2p-1}_{E}(\mathcal{Y}^{an}_{\mathbb{C},C},\mathbb{Q})$, where $C\subset\bar{T}_{\mathbb{C}}^{o}$ is a smooth transversal slice of $S_{\mathbb{C}}$ at $t$ in particular $\mathop{\rm dim}\nolimits(C)=1$, $C\cap S_{\mathbb{C}}=\left\{t\right\}$ and $\mathcal{Y}_{C}$ is smooth, and
    $\mathbb{V}^{p}_{S}(\tilde{\lambda})=c_{\mathcal{E}}(\mathbb{V}^{p}_{S}(\tilde{\lambda})),\;\mathbb{V}^{p}_{S}(\tilde{\lambda})\subset E_{DR,\mathcal{E}_{\mathbb{C}}}^{2p-1}(\mathcal{Y}_{\bar{T}^{o},\mathbb{C}}/\bar{T}_{\mathbb{C}}^{o}),$
    with
    $c_{\mathcal{E}}:(E_{DR,\mathcal{E}_{\mathbb{C}}}^{2p-1}(\mathcal{Y}_{\bar{T}^{o},\mathbb{C}}/\bar{T}_{\mathbb{C}}^{o}),F,W)\to(E^{2p}_{DR}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}}),F,W).$
    We have for each $t^{\prime}\in S(\mathbb{C})$, the isomorphism
    $\oplus_{cardI=2}F^{p}H^{2p}(\mathcal{E}^{an}_{I\mathbb{C},t^{\prime}},\mathbb{Q})=F^{p}\mathop{\rm Gr}\nolimits^{W}_{2p}H^{2p-2}((\bar{\mathcal{Y}_{\mathbb{C},C^{\prime}}}\backslash\mathcal{E}_{t^{\prime}})^{an},\mathbb{Q})$
    $\xrightarrow{c(\mathbb{Z}(\mathcal{Y}_{\mathbb{C},C^{\prime}}\backslash\mathcal{E}_{t^{\prime}}))}F^{p}\mathop{\rm Gr}\nolimits^{W}_{2p}H^{2p-1}_{E}(\bar{\mathcal{Y}^{an}_{\mathbb{C},C^{\prime}}},\mathbb{Q}),$
    where $C^{\prime}\subset\bar{T}_{\mathbb{C}}^{o}$ is a smooth transversal slice of $S$ at $t^{\prime}$ (hence $\mathcal{Y}_{C^{\prime}}$ is smooth) and $\bar{\mathcal{Y}_{C^{\prime}}}$ is a smooth compactification of $\mathcal{Y}_{C^{\prime}}$. Hence, assuming the Hodge conjecture for smooth projective varieties of dimension less or equal to $d-1$,
    $\mathbb{V}^{p}_{S}(\lambda)\subset\mathop{\rm Gr}\nolimits^{W}_{2p}E_{DR,\mathcal{E}_{\mathbb{C}}}^{2p-1}(\mathcal{Y}_{\bar{T}^{o},\mathbb{C}}/\bar{T}_{\mathbb{C}}^{o}),\;\mathbb{V}^{p}_{S}(\lambda)=c_{\mathcal{E}}(\mathbb{V}^{p}_{S}(\lambda))\subset E^{2p}_{DR}(\mathcal{E}_{0\mathbb{C}}/S_{\mathbb{C}})\subset\mathop{\rm Gr}\nolimits^{W}_{2p}E^{2p}_{DR}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})$
    are defined over $\bar{\mathbb{Q}}$ and its Galois conjugates are also components of the locus of Hodge classes, since $\mathop{\rm dim}\nolimits(\mathcal{E}_{I\mathbb{C},t^{\prime}})=d-1$ for $I\subset[0,\cdots,r]$ such that $cardI=2$ and $t^{\prime}\in S(\mathbb{C})$.
  • Consider now the case where $Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda})\neq 0$. By [5] theorem 1, the Hodge conjecture holds for projective hypersurfaces, hence
    $\mathbb{V}_{S}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))=\overline{\mathbb{V}_{T^{o}}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))}\cap p^{-1}(S_{\mathbb{C}})\subset E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T_{\mathbb{C}}^{o}),$
    where $p:=p\otimes\mathbb{C}:E^{2p}_{DR,V_{0}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T^{o}_{\mathbb{C}})\to\bar{T}^{o}_{\mathbb{C}}$ is the projection, is an algebraic subvariety defined over $\bar{\mathbb{Q}}$ and its Galois conjugates are also components of the locus of Hodge classes, where equality follows from [6] lemma 2.11. Hence,
    $\mathbb{V}_{S}^{p}(\lambda)=q(Sp_{\mathcal{E}/\tilde{f}_{X}}^{-1}(\mathbb{V}_{S}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))))\subset E^{2p}_{DR}(\mathcal{E}_{0\mathbb{C}}/S_{\mathbb{C}})$
    is an algebraic subvariety defined over $\bar{\mathbb{Q}}$ and its Galois conjugates are also components of the locus of Hodge classes, where the second equality follows from the fact that $Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda})\neq 0$ and (1).
The fact that $\mathbb{V}_{S}^{p}(\lambda)\subset E^{2p}_{DR}(\mathcal{E}_{0\mathbb{C}}/S_{\mathbb{C}})$ is an algebraic subvariety defined over $\bar{\mathbb{Q}}$ and its Galois conjugates are also components of the locus of Hodge classes gives, by [4] theorem 4,
$\lambda=[Z]\in H^{2p}(E^{an}_{0},\mathbb{Q}),\;Z\in\mathcal{Z}^{p}(E_{0}).$
We recall the key point is that since $\mathbb{V}_{S}^{p}(\lambda)\subset E^{2p}_{DR}(\mathcal{E}_{0\mathbb{C}}/S_{\mathbb{C}})$ is an algebraic subvariety defined over $\bar{\mathbb{Q}}$ and its Galois conjugate are also components of the locus of Hodge classes, we have for each $\theta\in Aut(\mathbb{C}/\mathbb{Q})$,
$\theta(\lambda)\in Gal(\bar{\mathbb{Q}}/\mathbb{Q})(\mathbb{V}_{S}^{p}(\lambda)_{\theta})\subset E^{2p}_{DR}(\mathcal{E}_{0\mathbb{C},\theta}/S_{\mathbb{C},\theta}),$
which implies that $\theta(\lambda)\in H^{2p}(E^{an}_{0,\theta},\mathbb{Q})\subset H^{2p}(E^{an}_{0,\theta},\mathbb{C})$, that is $\lambda\in H^{2p}(E^{an}_{0},\mathbb{Q})$ is absolute Hodge and we apply [4] corollary 1. Since $p\in\mathbb{Z}$ and $\lambda\in F^{p}H^{2p}(E_{0}^{an},\mathbb{Q})$ are arbitrary, this proves the Hodge conjecture for $E_{0}$. Since $X$ is birational to $E_{0}$, the Hodge conjecture for $E_{0}$ and the induction hypothesis implies the Hodge conjecture for $X$.
I am grateful to professor F.Mokrane for help and support during this work.
2 Preliminaries and notations
2.1 Notations
  • Denote by $\mathop{\rm Top}\nolimits$ the category of topological spaces and $\mathop{\rm RTop}\nolimits$ the category of ringed spaces.
  • Denote by $\mathop{\rm Cat}\nolimits$ the category of small categories and $\mathop{\rm RCat}\nolimits$ the category of ringed topos.
  • For $\mathcal{S}\in\mathop{\rm Cat}\nolimits$ and $X\in\mathcal{S}$, we denote $\mathcal{S}/X\in\mathop{\rm Cat}\nolimits$ the category whose objects are $Y/X:=(Y,f)$ with $Y\in\mathcal{S}$ and $f:Y\to X$ is a morphism in $\mathcal{S}$, and whose morphisms $\mathop{\rm Hom}\nolimits((Y^{\prime},f^{\prime}),(Y,f))$ consists of $g:Y^{\prime}\to Y$ in $\mathcal{S}$ such that $f\circ g=f^{\prime}$.
  • Let $(\mathcal{S},O_{S})\in\mathop{\rm RCat}\nolimits$ a ringed topos with topology $\tau$. For $F\in C_{O_{S}}(\mathcal{S})$, we denote by $k:F\to E_{\tau}(F)$ the canonical flasque resolution in $C_{O_{S}}(\mathcal{S})$ (see [3]). In particular for $X\in\mathcal{S}$, $H^{*}(X,E_{\tau}(F))\xrightarrow{\sim}\mathbb{H}_{\tau}^{*}(X,F)$.
  • For $f:\mathcal{S}^{\prime}\to\mathcal{S}$ a morphism with $\mathcal{S},\mathcal{S}^{\prime}\in\mathop{\rm RCat}\nolimits$, endowed with topology $\tau$ and $\tau^{\prime}$ respectively, we denote for $F\in C_{O_{S}}(\mathcal{S})$ and each $j\in\mathbb{Z}$,
    • $f^{*}:=H^{j}\Gamma(\mathcal{S},k\circ\mathop{\rm ad}\nolimits(f^{*},f_{*})(F)):\mathbb{H}^{j}(\mathcal{S},F)\to\mathbb{H}^{j}(\mathcal{S}^{\prime},f^{*}F)$,
    • $f^{*}:=H^{j}\Gamma(\mathcal{S},k\circ\mathop{\rm ad}\nolimits(f^{*mod},f_{*})(F)):\mathbb{H}^{j}(\mathcal{S},F)\to\mathbb{H}^{j}(\mathcal{S}^{\prime},f^{*mod}F)$,
    the canonical maps.
  • For $m:A\to B$, $A,B\in C(\mathcal{A})$, $\mathcal{A}$ an additive category, we denote $c(A):\mathop{\rm Cone}\nolimits(m:A\to B)\to A[1]$ and $c(B):B\to\mathop{\rm Cone}\nolimits(m:A\to B)$ the canonical maps.
  • Denote by $\mathop{\rm Sch}\nolimits\subset\mathop{\rm RTop}\nolimits$ the subcategory of schemes (the morphisms are the morphisms of locally ringed spaces). We denote by $\mathop{\rm PSch}\nolimits\subset\mathop{\rm Sch}\nolimits$ the full subcategory of proper schemes. For a field $k$, we consider $\mathop{\rm Sch}\nolimits/k:=\mathop{\rm Sch}\nolimits/\mathop{\rm Spec}\nolimits k$ the category of schemes over $\mathop{\rm Spec}\nolimits k$. The objects are $X:=(X,a_{X})$ with $X\in\mathop{\rm Sch}\nolimits$ and $a_{X}:X\to\mathop{\rm Spec}\nolimits k$ a morphism and the morphisms are the morphisms of schemes $f:X^{\prime}\to X$ such that $f\circ a_{X^{\prime}}=a_{X}$. We then denote by
    • $\mathop{\rm Var}\nolimits(k)=\mathop{\rm Sch}\nolimits^{ft}/k\subset\mathop{\rm Sch}\nolimits/k$ the full subcategory consisting of algebraic varieties over $k$, i.e. schemes of finite type over $k$,
    • $\mathop{\rm PVar}\nolimits(k)\subset\mathop{\rm QPVar}\nolimits(k)\subset\mathop{\rm Var}\nolimits(k)$ the full subcategories consisting of quasi-projective varieties and projective varieties respectively,
    • $\mathop{\rm PSmVar}\nolimits(k)\subset\mathop{\rm SmVar}\nolimits(k)\subset\mathop{\rm Var}\nolimits(k)$, $\mathop{\rm PSmVar}\nolimits(k):=\mathop{\rm PVar}\nolimits(k)\cap\mathop{\rm SmVar}\nolimits(k)$, the full subcategories consisting of smooth varieties and smooth projective varieties respectively.
  • Denote by $\mathop{\rm AnSp}\nolimits(\mathbb{C})\subset\mathop{\rm RTop}\nolimits$ the subcategory of analytic spaces over $\mathbb{C}$, and by $\mathop{\rm AnSm}\nolimits(\mathbb{C})\subset\mathop{\rm AnSp}\nolimits(\mathbb{C})$ the full subcategory of smooth analytic spaces (i.e. complex analytic manifold).
  • For $X\in\mathop{\rm Var}\nolimits(k)$ and $X=\cup_{i\in I}X_{i}$ with $i_{i}:X_{i}\hookrightarrow X$ closed embeddings, we denote $X_{\bullet}\in\mathop{\rm Fun}\nolimits(\Delta,\mathop{\rm Var}\nolimits(k))$ the associated simplicial space, with for $J\subset I$, $i_{IJ}:X_{I}:=\cap_{i\in I}X_{i}\hookrightarrow X_{J}:=\cap_{i\in J}X_{i}$ the closed embedding.
  • Let $(X,O_{X})\in\mathop{\rm RTop}\nolimits$. We consider its De Rham complex $\Omega_{X}^{\bullet}:=DR(X)(O_{X})$.
    • Let $X\in\mathop{\rm Sch}\nolimits$. Considering its De Rham complex $\Omega_{X}^{\bullet}:=DR(X)(O_{X})$, we have for $j\in\mathbb{Z}$ its De Rham cohomology $H^{j}_{DR}(X):=\mathbb{H}^{j}(X,\Omega^{\bullet}_{X})$.
    • Let $X\in\mathop{\rm Var}\nolimits(k)$. Considering its De Rham complex $\Omega_{X}^{\bullet}:=\Omega_{X/k}^{\bullet}:=DR(X/k)(O_{X})$, we have for $j\in\mathbb{Z}$ its De Rham cohomology $H^{j}_{DR}(X):=\mathbb{H}^{j}(X,\Omega^{\bullet}_{X})$. The differentials of $\Omega_{X}^{\bullet}:=\Omega_{X/k}^{\bullet}$ are by definition $k$-linear, thus $H^{j}_{DR}(X):=\mathbb{H}^{j}(X,\Omega^{\bullet}_{X})$ has a structure of a $k$ vector space.
    • Let $X\in\mathop{\rm AnSp}\nolimits(\mathbb{C})$. Considering its De Rham complex $\Omega_{X}^{\bullet}:=DR(X)(O_{X})$, we have for $j\in\mathbb{Z}$ its De Rham cohomology $H^{j}_{DR}(X):=\mathbb{H}^{j}(X,\Omega^{\bullet}_{X})$.
  • For $Y\in\mathop{\rm Var}\nolimits(\mathbb{Q})$ and $X=V(I)=V(f_{1},\ldots,f_{r})\subset Y_{\mathbb{C}}$, we consider $\mathcal{X}=V(\tilde{I})=V(\tilde{f}_{1},\ldots,\tilde{f}_{r})\subset Y\times\mathbb{A}_{\mathbb{Q}}^{s}$ the canonical rational deformation, where
    $\mbox{if}\,f=\sum_{I\subset[1,\ldots,d]^{n}}a_{I}x_{1}^{n_{1}}\cdots x_{n}^{n_{n}}\in\mathbb{C}[x_{1},\ldots,x_{n}],$
    $\tilde{f}=\sum_{I\subset[1,\ldots,d]^{n}}a_{I}x_{1}^{n_{1}}\cdots x_{n}^{n_{n}}\in\mathbb{Q}[(a_{I})_{I\subset[1,\ldots,d]^{n}},x_{1},\ldots,x_{n}].$
  • For $X\in\mathop{\rm AnSp}\nolimits(\mathbb{C})$, we denote $\alpha(X):\mathbb{C}_{X}\hookrightarrow\Omega_{X}^{\bullet}$ the embedding in $C(X)$. For $X\in\mathop{\rm AnSm}\nolimits(\mathbb{C})$, $\alpha(X):\mathbb{C}_{X}\hookrightarrow\Omega_{X}^{\bullet}$ is an equivalence usu local by Poincare lemma.
  • We denote $\mathbb{I}^{n}:=[0,1]^{n}\in\mathop{\rm Diff}\nolimits(\mathbb{R})$ (with boundary). For $X\in\mathop{\rm Top}\nolimits$ and $R$ a ring, we consider its singular cochain complex
    $C^{*}_{\mathop{\rm sing}\nolimits}(X,R):=(\mathbb{Z}\mathop{\rm Hom}\nolimits_{\mathop{\rm Top}\nolimits}(\mathbb{I}^{*},X)^{\vee})\otimes R$
    and for $l\in\mathbb{Z}$ its singular cohomology $H^{l}_{\mathop{\rm sing}\nolimits}(X,R):=H^{n}C^{*}_{\mathop{\rm sing}\nolimits}(X,R)$. For $f:X^{\prime}\to X$ a continuous map with $X,X^{\prime}\in\mathop{\rm Top}\nolimits$, we have the canonical map of complexes
    $f^{*}:C^{*}_{\mathop{\rm sing}\nolimits}(X,R)\to C^{*}_{\mathop{\rm sing}\nolimits}(X,R),\sigma\mapsto f^{*}\sigma:=(\gamma\mapsto\sigma(f\circ\gamma)).$
    In particular, we get by functoriality the complex
    $C^{*}_{X,R\mathop{\rm sing}\nolimits}\in C_{R}(X),\;(U\subset X)\mapsto C^{*}_{\mathop{\rm sing}\nolimits}(U,R)$
    We recall that
    • For $X\in\mathop{\rm Top}\nolimits$ locally contractible, e.g. $X\in\mathop{\rm CW}\nolimits$, and $R$ a ring, the inclusion in $C_{R}(X)$ $c_{X}:R_{X}\to C^{*}_{X,R\mathop{\rm sing}\nolimits}$ is by definition an equivalence top local and that we get by the small chain theorem, for all $l\in\mathbb{Z}$, an isomorphism $H^{l}c_{X}:H^{l}(X,R_{X})\xrightarrow{\sim}H^{l}_{\mathop{\rm sing}\nolimits}(X,R)$.
    • For $X\in\mathop{\rm Diff}\nolimits(\mathbb{R})$, the restriction map
      $r_{X}:\mathbb{Z}\mathop{\rm Hom}\nolimits_{\mathop{\rm Diff}\nolimits(\mathbb{R})}(\mathbb{I}^{*},X)^{\vee}\to C^{*}_{\mathop{\rm sing}\nolimits}(X,R),\;w\mapsto w:(\phi\mapsto w(\phi))$
      is a quasi-isomorphism by Whitney approximation theorem.
  • Let $S\in\mathop{\rm AnSm}\nolimits(\mathbb{C})$ and $L\in\mathop{\rm Shv}\nolimits(S)$ a local system. Consider $E:=L\otimes O_{S}\in\mathop{\rm Vect}\nolimits_{\mathcal{D}}(S)$ the corresponding holomorphic vector bundle with integrable connection $\nabla$. Let $\pi_{S}:\tilde{S}\to S$ be the universal covering. Consider the canonical fiber
    $L_{0}:=\Gamma(\tilde{S},\pi_{S}^{*}L)=\Gamma(\tilde{S},\pi_{S}^{*mod}E)^{\nabla}.$
    For $\lambda\in L_{0}$, we will consider the flat leaf
    $\mathbb{V}_{S}(\lambda):=\pi_{S}(\lambda\times\tilde{S})\subset E,\;\pi_{S}:L_{0}\times\tilde{S}\to E,\,\pi_{S}(\nu,z):=(\nu(z),\pi_{S}(z)).$
    Note that for $t\in S$, $\mathbb{V}_{S}(\lambda)\cap p_{S}^{-1}(t)=\pi_{1}(S,t)(\lambda)$ is the orbit of $\lambda$ under the monodromy action, where $p_{S}:E\to S$ is the projection.
2.2 Birational projective varieties and Hodge conjecture
In this subsection, we recall that assuming the Hodge conjecture for smooth projective varieties of dimension less or equal to $d-1$, the property of satisfying the Hodge conjecture for connected smooth projective varieties of dimension $d$ is a birational invariant :
Proposition 1.
  • (i)
    Let $X\in\mathop{\rm PSmVar}\nolimits(\mathbb{C})$ connected of dimension $d$. Let $\epsilon:\tilde{X}_{Z}\to X$ be the blow up of $X$ along a smooth closed subvariety $Z\subset X$. Assume the Hodge conjecture hold for smooth complex projective varieties of dimension less or equal to $d-1$. Then the Hodge conjecture hold for $X$ if and only if it hold for $\tilde{X}_{Z}$.
  • (ii)
    Let $X,X^{\prime}\in\mathop{\rm PSmVar}\nolimits(\mathbb{C})$ connected of dimension $d$. Assume the Hodge conjecture hold for smooth complex projective varieties of dimension less or equal to $d-1$. If $X$ is birational to $X^{\prime}$, then the Hodge conjecture hold for $X$ if and only if it hold for $X^{\prime}$
Proof.
(i): Follows from the fact that
$(\epsilon^{*},\oplus_{l=1}^{c}i_{E*}((-).h^{l-1})):H^{k}(X,\mathbb{Q})\oplus\oplus_{l=1}^{c}H^{k-2l}(E,\mathbb{Q})\to H^{k}(\tilde{X}_{Z},\mathbb{Q})$
is an isomorphism of Hodge structures for each $k\in\mathbb{Z}$, where $c=\mathop{\rm codim}\nolimits(Z,X)$, $i_{E}:E\hookrightarrow\tilde{X}_{Z}$ is the closed embedding, $\epsilon_{|E}:E\to Z$ being a projective vector bundle.
(ii):Follows from (i) since if $\pi:X^{o}\to X^{\prime}$ is a birational map, $X^{o}\subset X$ being an open subset, then $X$ is connected to $X^{\prime}$ by a sequence of blow up of smooth projective varieties with smooth center by [1]. ∎
2.3 A family of smooth projective hypersurfaces associated to a smooth projective variety
In this subsection, we recall that given a smooth complex projective variety $X$, there exists a family of smooth complex projective hypersurfaces which degenerates into a normal crossing divisor with one irreducible component birational to $X$.
Proposition 2.
Let $X\in\mathop{\rm PSmVar}\nolimits(\mathbb{C})$ connected of dimension $d$. Then there exists a family of smooth projective hypersurfaces $f_{X}:Y\to\mathbb{A}^{1}_{\mathbb{C}}$ with
  • $Y\in\mathop{\rm SmVar}\nolimits(\mathbb{C})$, $f_{X}$ flat projective,
  • $Y_{s}:=f_{X}^{-1}(s)\subset\mathbb{P}^{d+1}_{\mathbb{C}}$ for $s\in\mathbb{A}^{1}_{\mathbb{C}}\backslash 0$ are smooth projective hypersurfaces,
  • $E:=f_{X}^{-1}(0)=\cup_{i=0}^{r}E_{i}=\epsilon^{-1}(Y_{0})\subset Y$ is a normal crossing divisor, where $Y_{0}\subset\mathbb{P}^{d+1}_{\mathbb{C}}$ is a possibly singular hypersurface birationnal to $X$, $\epsilon:(Y,E)\to(Y^{\prime},Y_{0})$ a desingularization, and $E_{0}$ is birational to $X$,
Proof.
See [2] lemma 1.5.8 . It use the fact that $X$ is birational to a possibly singular projective hypersurface $Y_{0}$ and consider the desingularization of a generic pencil $f^{\prime}_{X}:Y^{\prime}=V(f)\subset\mathbb{P}^{d+1}_{\mathbb{C}}\times\mathbb{A}^{1}_{\mathbb{C}}\to\mathbb{A}^{1}_{\mathbb{C}}$ passing through $Y_{0}$ and $f_{X}=f^{\prime}_{X}\circ\epsilon$. ∎
2.4 The nearby cycle functor and the specialization map
Recall from [3] that for $k\subset\mathbb{C}$ a subfield and $S\in\mathop{\rm Var}\nolimits(k)$ quasi-projective, and $l:S\hookrightarrow\tilde{S}$ a closed embedding with $\tilde{S}\in\mathop{\rm SmVar}\nolimits(k)$, we have the full subcategory
$\iota:MHM_{k,gm}(S)\hookrightarrow\mathop{\rm PSh}\nolimits_{\mathcal{D}fil,S}(\tilde{S})\times_{I}P(S^{an}_{\mathbb{C}})$
consisting of geometric mixed Hodge module whose De Rham part is defined over $k$, where, using the fact that the $V$-filtration is defined over $k$, $\mathop{\rm PSh}\nolimits_{\mathcal{D}(1,0)fil,S}(\tilde{S})\times_{I}P_{fil}(S^{an}_{\mathbb{C}})$ is the category
  • whose objects are $((M,F,W),(K,W),\alpha)$ where $(M,F,W)$ is a filtered $D_{\tilde{S}}$ module supported on $S$, where $l:S\hookrightarrow\tilde{S}$ is a closed embedding and $\tilde{S}\in\mathop{\rm SmVar}\nolimits(k)$, $(K,W)$ is a filtered Perverse sheaf on $S^{an}_{\mathbb{C}}$, and $\alpha:l_{*}(K,W)\otimes\mathbb{C}\to DR(\tilde{S})((M,W)^{an})$ is an isomorphism in $D_{fil}(\tilde{S}^{an}_{\mathbb{C}})$.
  • whose morphism between $((M,F,W),(K,W),\alpha)$ and $((M^{\prime},F,W),(K^{\prime},W),\alpha^{\prime})$, are couples of morphisms $(\phi:(M,F,W)\to(M^{\prime},F,W),\psi:(K,W)\to(K^{\prime},W))$ such that $\phi\circ\alpha=\alpha^{\prime}\circ\psi$.
Then, by [3], we get for $S\in\mathop{\rm Var}\nolimits(k)$ quasi-projective and $l:S\hookrightarrow\tilde{S}$ is a closed embedding with $\tilde{S}\in\mathop{\rm SmVar}\nolimits(k)$ an embedding
$\iota:D(MHM_{k,gm}(S))\hookrightarrow D_{\mathcal{D}fil,S}(\tilde{S})\times_{I}D(S^{an}_{\mathbb{C}})$
consisting of mixed Hodge module whose De Rham part is defined over $k$, where $D_{\mathcal{D}(1,0)fil,S}(\tilde{S})\times_{I}D_{fil,c}(S^{an}_{\mathbb{C}})$ is the category
  • whose objects are $((M,F,W),(K,W),\alpha)$ where $(M,F,W)$ is a complex of filtered $D_{\tilde{S}}$ module supported on $S$, $(K,W)$ is a filtered complex of presheaves on $S^{an}_{\mathbb{C}}$ whose cohomology are constructible sheaves, and $\alpha:l_{*}(K,W)\otimes\mathbb{C}\to DR(\tilde{S})((M,W)^{an})$ is an isomorphism in $D_{fil}(\tilde{S}^{an}_{\mathbb{C}})$.
  • whose morphism are given in [3],
We have then the six functors formalism :
$D(MHM_{k,gm}(-)):\mathop{\rm Var}\nolimits(k)\mapsto\mathop{\rm TriCat}\nolimits,\;S\mapsto D(MHM_{k,gm}(S)),\;(f:T\to S)\mapsto$
$(f^{*},f_{*}):D(MHM_{k,gm}(S))\to D(MHM_{k,gm}(T)),\;(f_{!},f^{!}):D(MHM_{k,gm}(T))\to D(MHM_{k,gm}(S))$
and for $D\subset S$ a (Cartier) divisor the nearby cycle functor
$\psi_{D}:D(MHM_{k,gm}(S))\to D(MHM_{k,gm}(D)),$
$\psi_{D}((M,F,W),(K,W),\alpha):=(\psi_{D}(M,F,W),\psi_{D}(K,W),\psi_{D}(\alpha)).$
For $T\in\mathop{\rm Var}\nolimits(k)$ and $\epsilon_{\bullet}:T_{\bullet}\to T$ a desingularization with $T_{\bullet}\in\mathop{\rm Fun}\nolimits(\Delta,\mathop{\rm SmVar}\nolimits(k))$, we have
$\mathbb{Z}_{T}^{Hdg}:=a_{T}^{*Hdg}\mathbb{Z}^{Hdg}=(\epsilon_{\bullet*}E_{zar}(O_{T_{\bullet}},F,W),\epsilon_{\bullet*}E_{usu}(\mathbb{Z}_{T^{an}_{\bullet}},W),\alpha(T_{\bullet}))$
For $f:X\to S$ a projective morphism with $X,S\in\mathop{\rm SmVar}\nolimits(k)$, we denote
$E_{Hdg}(X/S):=f_{*}\mathbb{Z}_{X}^{Hdg}:=(\int_{f}(O_{X},F_{b}),Rf_{*}\mathbb{Z}_{X^{an}_{\mathbb{C}}},f_{*}\alpha(X))\in D(MHM_{k,gm}(S))$
and $E^{j}_{Hdg}(X/S):=H^{j}E_{Hdg}(X/S)$ for $j\in\mathbb{Z}$. If moreover $Z\subset X$ is a closed subset, we denote
$E_{Hdg,Z}(X/S):=f_{*}\Gamma_{Z}^{Hdg}\mathbb{Z}_{X}^{Hdg}:=(\int_{f}\Gamma_{Z}^{Hdg}(O_{X},F_{b}),Rf_{*}R\Gamma_{Z}\mathbb{Z}_{X^{an}_{\mathbb{C}}},f_{*}\Gamma_{Z}\alpha(X))\in D(MHM_{k,gm}(S))$
[…]
$E_{Hdg}(X/S):=f_{*}\mathbb{Z}_{X}^{Hdg}:=(\int_{p}\Gamma^{\vee,Hdg}_{X}(O_{\mathbb{P}^{N}\times S},F),Rf_{*}\mathbb{Z}_{X^{an}_{\mathbb{C}}},f_{*}\alpha(X))\in D(MHM_{k,gm}(S))$
and $E^{j}_{Hdg}(X/S):=H^{j}E_{Hdg}(X/S)$ for $j\in\mathbb{Z}$. In the particular case where $E_{Hdg}^{j}(X/S)$ is a variation of mixed Hodge structure, the first factor of
$E^{j}_{Hdg}(X/S):=(E^{j}_{DR}(X/S),R^{j}f_{*}\mathbb{Z}_{X^{an}_{\mathbb{C}}},H^{j}f_{*}\alpha(X))\in MHM_{k,gm}(S)$
is a vector bundle.
We then have the specialization map :
Definition-Proposition 1.
Let $k\subset\mathbb{C}$ be a subfield.
  • (i)
    Let $f^{\prime}:X^{\prime}\to T$ be a projective morphism with $X^{\prime},T\in\mathop{\rm SmVar}\nolimits(k)$. Assume that
    • $f^{o}:=f\otimes_{T}T:X^{o}\to T^{o}$ is smooth, $T^{o}\subset T$ is an open subset,
    • $X^{\prime}_{S}:=f^{-1}(S)\subset X^{\prime}$, $S:=T\backslash T^{o}$. We have by definition $X^{o}=X^{\prime}\backslash X^{\prime}_{S}$.
    We have then, see [3], the distinguished triangles in $D(MHM_{k,gm}(X_{S}))$,
    $\mathbb{Z}_{X^{\prime}_{S}}^{Hdg}:=i_{X^{\prime}_{S}}^{*}\mathbb{Z}_{X^{\prime}}^{Hdg}\xrightarrow{Sp_{X^{\prime}_{S}/X^{\prime}}}\psi_{X^{\prime}_{S}}\mathbb{Z}^{Hdg}_{X^{o}}:=(\psi_{X^{\prime}_{S}}(O_{X^{\prime}},F),\psi_{X^{\prime}_{S}}\mathbb{Z}_{X^{o}},\psi_{X^{\prime}_{S}}\alpha(X^{\prime}))$
    $\xrightarrow{c(\psi_{X^{\prime}_{S}}\mathbb{Z}^{Hdg}_{X^{o}})}\phi_{X^{\prime}_{S}}\mathbb{Z}^{Hdg}_{X^{o}}:=\mathop{\rm Cone}\nolimits(Sp_{X^{\prime}/X_{S}})\xrightarrow{c(\mathbb{Z}_{X^{\prime}_{S}}^{Hdg})}\mathbb{Z}_{X^{\prime}_{S}}^{Hdg}[1].$
    Applying the functor $(f^{\prime}\circ i_{X^{\prime}_{S}})_{*}$, we obtain the generalized distinguished triangle in $D(MHM_{k,gm}(S))$
    $E_{Hdg}(X^{\prime}_{S}/S)\xrightarrow{sp_{X^{\prime}_{S}/f}:=(f\circ i_{X^{\prime}_{S}})_{*}Sp_{X^{\prime}_{S}/X^{\prime}}}$
    $(f^{\prime}\circ i_{X^{\prime}_{S}})_{*}(\psi_{X^{\prime}_{S}}\mathbb{Z}^{Hdg}_{X^{o}})=\psi_{0}E_{Hdg}(X^{o}/T^{o})\xrightarrow{(f^{\prime}\circ i_{X^{\prime}_{S}})_{*}c(\Phi_{X_{0}}\mathbb{Z}^{Hdg}_{X^{o}})}$
    $(f^{\prime}\circ i_{X^{\prime}_{S}})_{*}(\phi_{X_{S}}\mathbb{Z}^{Hdg}_{X^{o}})=\phi_{S}E_{Hdg}(X^{o}/T^{o})\xrightarrow{(f^{\prime}\circ i_{X^{\prime}_{S}})_{*}c(\mathbb{Z}^{Hdg}_{X^{\prime}})}E_{Hdg}(X^{\prime}_{S}/S)[1]$
  • (ii)
    Let $f:X\to T$ be a projective morphism with $X,T\in\mathop{\rm SmVar}\nolimits(k)$. Assume that
    • $f^{o}:=f\otimes_{T}T^{o}:X^{o}\to T^{o}$ is smooth, $T^{o}\subset T$ an open subset
    • $S:=T\backslash T^{o}$, $E:=f^{-1}(S)=\cup_{i=0}^{r}E_{i}\subset X$ is a normal crossing divisor and that for $I\subset[0,\ldots r]$, $f_{0,I}:=f_{|E_{I}}:E_{I}:=\cap_{i\in I}E_{i}\to S$ are smooth. We have by definition $X^{o}=X\backslash E$.
    Denote for $I\subset J\subset[0,\ldots r]$, $i_{JI}:E_{J}\hookrightarrow E_{I}$ and $i_{I}:E_{I}\hookrightarrow E$, $i_{E}:E\hookrightarrow X$ the closed embeddings. In this particular case, c.f. [8], the distinguished triangle in $D(MHM_{k,gm}(E))$ given in (i) become
    $\mathbb{Z}_{E}^{Hdg}:=i_{E}^{*}\mathbb{Z}_{X}^{Hdg}=((i_{\bullet*}\Omega^{\bullet}_{E^{\bullet}},F,W),(\mathbb{Z}_{E_{\bullet}},W),\alpha(E_{\bullet}))\xrightarrow{Sp_{E/X}}$
    $\psi_{E}\mathbb{Z}^{Hdg}_{X^{o}}=((i_{E}^{*mod}\Omega^{\bullet}_{X/T}(\mathop{\rm log}\nolimits E),F,W),\psi_{E}\mathbb{Z}_{X^{o}},\psi_{E}\alpha(X))\xrightarrow{c(\psi_{E}\mathbb{Z}^{Hdg}_{X^{o}})}$
    $\phi_{E}\mathbb{Z}^{Hdg}_{X^{o}}:=\mathop{\rm Cone}\nolimits(Sp_{E/X})\xrightarrow{c(\mathbb{Z}_{E}^{Hdg})}\mathbb{Z}_{E}^{Hdg}[1].$
    Recall the map $Sp_{E/X}$ is the defined as the factorization
    $\mathbb{Z}_{X}^{Hdg}\xrightarrow{\mathop{\rm ad}\nolimits(i_{E}^{*},i_{E*})(\mathbb{Z}_{X}^{Hdg}}\mathbb{Z}_{E}^{Hdg}\xrightarrow{Sp_{E/X}}\psi_{E}\mathbb{Z}^{Hdg}_{X^{o}}.$
    Applying the functor $(f\circ i_{E})_{*}$, we obtain the generalized distinguished triangle in $D(MHM_{k,gm}(S))$
    $E_{Hdg}(E/S)\xrightarrow{Sp_{E/f}:=(f\circ i_{E})_{*}Sp_{E/X}}(f\circ i_{E})_{*}(\psi_{E}\mathbb{Z}^{Hdg}_{X^{o}})=\psi_{S}E_{Hdg}(X^{o}/T^{o})$
    $\xrightarrow{(f\circ i_{E})_{*}c(\psi_{E}\mathbb{Z}^{Hdg}_{X^{o}})}(f\circ i_{E})_{*}(\phi_{E}\mathbb{Z}^{Hdg}_{X^{o}})=\phi_{S}E_{Hdg}(X^{o}/T^{o})\xrightarrow{(f\circ i_{E})_{*}c(\mathbb{Z}^{Hdg}_{E})}E_{Hdg}(E/S)[1].$
    If $f=f^{\prime}\circ\epsilon$ where $\epsilon:(X,E)\to(X,X_{S})$ is a desingularization, we have $Sp_{X_{S}/f^{\prime}}=Sp_{E/f}\circ\epsilon^{*}$.
Proof.
See [8]. ∎
Lemma 1.
Let $k\subset\mathbb{C}$ be a subfield. Let $f:X\to T$ be a projective morphism with $X,T\in\mathop{\rm SmVar}\nolimits(k)$. Assume that
  • $f^{o}:=f\otimes_{T}T^{o}:X^{o}\to T^{o}$ is smooth of relative dimension $n$, $T^{o}\subset T$ an open subset
  • $S:=T\backslash T^{o}=V(s)\subset T$ is a divisor, $E:=f^{-1}(S)=\cup_{i=0}^{r}E_{i}\subset X$ is a normal crossing divisor and that for $I\subset[0,\ldots r]$, $f_{0,I}:=f_{|E_{I}}:E_{I}:=\cap_{i\in I}E_{i}\to S$ are smooth. We have by definition $X^{o}=X\backslash E$.
Denote $j_{0}:E^{o}_{0}:=E_{0}\backslash\cup_{i=1}^{r}E_{i}\xrightarrow{j^{\prime}_{0}}E_{0}\xrightarrow{i_{0}}E$ the embedding.
  • (i)
    We have, for each $k\in\mathbb{Z}$, a canonical map in $MHM_{k,gm}(S)$
    $c_{E}:E^{k-1}_{Hdg,E}(X/T)\xrightarrow{c_{V}}\phi_{S}E^{k-1}_{Hdg}(X^{o}/T^{o})\xrightarrow{H^{k-1}(f\circ i_{E})_{*}c(\mathbb{Q}^{Hdg}_{E})}E^{k}_{Hdg}(E/S).$
  • (ii)
    If $H^{2p-1}(X_{t},\mathbb{Q})=H^{2p-2}(X_{t},\mathbb{Q})=0$ for $t\in T\backslash S$, then
    $c_{E}:E^{k-1}_{Hdg,E}(X/T)\xrightarrow{c_{V}}\phi_{S}E^{k-1}_{Hdg}(X^{o}/T^{o})$
    $\xrightarrow{H^{k-1}(f\circ i_{E})_{*}c(\mathbb{Q}^{Hdg}_{E})}\ker(sp_{E/f}:E^{k}_{Hdg}(E/S)\to\psi_{S}E^{k}_{Hdg}(X^{o}/T^{o})).$
    is an isomorphism.
Proof.
(i):See [8] for the definition of the map $c_{V}$. Recall that $Var=DR(S)(s)$.
(ii):Follows from the exact sequences in $MHM_{k,gm}(S)$
[…]
$\cdots\to\psi_{S}E_{Hdg}^{k-1}(X^{o}/T^{o})\xrightarrow{H^{k-1}(f\circ i_{E})_{*}c(\psi_{E}\mathbb{Q}^{Hdg}_{X^{o}})}$
$\phi_{S}E_{Hdg}^{k-1}(X^{o}/T^{o})\xrightarrow{H^{k-1}(f\circ i_{E})_{*}c(\mathbb{Q}^{Hdg}_{E})}E_{Hdg}^{k}(E/S)\xrightarrow{Sp_{E/f}}\psi_{S}E_{Hdg}^{k}(X^{o}/T^{o})\to\cdots,$
given in [8] for the first one. ∎
3 Hodge conjecture for smooth projective varieties
Theorem 1.
Let $X\in\mathop{\rm PSmVar}\nolimits(\mathbb{C})$. Then Hodge conjecture hold for $X$. That is, if $p\in\mathbb{Z}$ and $\lambda\in F^{p}H^{2p}(X^{an},\mathbb{Q})$, $\lambda=[Z]$ with $Z\in\mathcal{Z}^{p}(X)$.
Proof.
The Hodge conjecture is true for curves and surfaces. Assume that the Hodge conjecture is true for smooth projective varieties of dimension less or equal to $d-1$. Let $X\in\mathop{\rm PSmVar}\nolimits(\mathbb{C})$ of dimension $d$. Up to split $X$ into its connected components, we may assume that $X$ is connected of dimension $d$. Then $X$ is birational to a possibly singular hypersurface $Y_{0_{X}}\subset\mathbb{P}^{d+1}_{\mathbb{C}}$ of degree $l_{X}:=deg(\mathbb{C}(X)/\mathbb{C}(x_{1},\cdots,x_{d}))$. Consider
$f:\mathcal{Y}^{\prime}_{\mathbb{C}}=V(\tilde{f})\subset\mathbb{P}^{d+1}_{\mathbb{C}}\times\mathbb{P}(l_{X})\to\mathbb{P}(l_{X})$
the universal family of projective hypersurfaces of degree $l_{X}$, and $0_{X}\in\mathbb{P}(l_{X})$ the point corresponding to $Y_{0_{X}}$. Denote $\Delta\subset\mathbb{P}(l_{X})$ the discriminant locus parametrizing the hypersurfaces which are singular. Let $\bar{S}:=\bar{0_{X}}^{\bar{\mathbb{Q}}}\subset\mathbb{P}(l_{X})$ be the $\bar{\mathbb{Q}}$-Zariski closure of $0_{X}$ inside $\mathbb{P}(l_{X})$. If $Y_{0_{X}}$ is smooth, then $X$ satisfy the hodge conjecture since $Y_{0_{X}}$ satisfy the Hodge conjecture ([5]) and by proposition 1(ii) and the induction hypothesis since $X$ is birational to $Y_{0_{X}}$. Hence, we may assume that $Y_{0_{X}}$ is singular, that is $0_{X}\in\Delta$. Then, $\bar{S}\subset\Delta$ since $\Delta\subset\mathbb{P}(l_{X})$ is defined over $\bar{\mathbb{Q}}$. Let $\bar{T}\subset\mathbb{P}(l_{X})$ be an irreducible closed subvariety of dimension $\mathop{\rm dim}\nolimits(\bar{S})+1$ defined over $\bar{\mathbb{Q}}$ such that
$\bar{S}\subset\bar{T}\;\mbox{and}\;T:=\bar{T}\cap(\mathbb{P}(l_{X})\backslash\Delta)\neq\emptyset.$
We take for $\bar{T}=V(f_{1},\cdots,f_{s})\subset\mathbb{P}(l_{X})$ an irreducible complete intersection defined over $\bar{\mathbb{Q}}$ which intersect $\Delta$ properly and such that
  • $V(f_{1})\supset\mathop{\rm sing}\nolimits(\Delta)\cup\bar{S}$
  • for each $r\in[1,\cdots,s]$, $V(f_{1},\cdots,f_{r})\supset\mathop{\rm sing}\nolimits(V(f_{1},\cdots,f_{r-1})\cap\Delta)\cup\bar{S}$ contain inductively the singular locus of the intersection with $\Delta$ and $\bar{S}$.
Then by the weak Lefchetz hyperplane theorem for homotopy groups, $T\subset\mathbb{P}(l_{X})\backslash\Delta$ is not contained in a weakly special subvariety since
$i_{T*}:\pi_{1}(T_{\mathbb{C}})\to\pi_{1}(\mathbb{P}(l_{X})\backslash\Delta)$
is an isomorphism if $\mathop{\rm dim}\nolimits\bar{T}\geq 2$ and surjective if $\mathop{\rm dim}\nolimits\bar{T}=1$. We have then $\bar{T}\cap\Delta=\bar{S}\cup\bar{S}^{\prime}$, where $\bar{S}^{\prime}\subset\bar{T}$ is a divisor. We then consider the family of algebraic varieties defined over $\mathbb{Q}$
$\tilde{f}_{X}:\mathcal{Y}_{\bar{T}^{o}}\xrightarrow{\epsilon}\mathcal{Y}^{\prime}_{\bar{T}^{o}}\xrightarrow{f\times_{\mathbb{P}(l_{X})}\bar{T}^{o}}\bar{T}^{o}$
where
  • $\tilde{f}_{X}:\mathcal{Y}\xrightarrow{\epsilon}\mathcal{Y}^{\prime}_{\bar{T}}\xrightarrow{f\times_{\mathbb{P}(l_{X})}\bar{T}}\bar{T}$ with $\epsilon:(\mathcal{Y},\mathcal{E}^{\prime})\to(\mathcal{Y}^{\prime}_{\bar{T}},\mathcal{Y}^{\prime}_{\bar{S}}\cup\mathcal{Y}^{\prime}_{\bar{S}^{\prime}})$ is a desingularization over $\mathbb{Q}$, in particular $\mathcal{E}^{\prime}=\mathcal{E}\cup\mathcal{E}^{\prime\prime}$ with $\mathcal{E}=\cup_{i=0}^{r}\mathcal{E}_{i}:=\tilde{f}_{X}^{-1}(\bar{S})\subset\mathcal{Y}$ which is a normal crossing divisor, in particular $\mathcal{Y}_{T}=\mathcal{Y}^{\prime}_{T}$,
  • $S\subset\bar{S}$ is the open subset such that $S$ is smooth and $\tilde{f}_{X}\times_{\bar{T}}S:\mathcal{E}_{I}:=\cap_{i\in I}\mathcal{E}_{i}\to S$ is smooth projective for each $I\subset[0,\ldots,r]$, $\bar{T}^{o}\subset\bar{T}$ is an open subset such that $\bar{T}^{o}\cap\Delta=S$ and $T^{o}:=\bar{T}^{o}\cap T$ is smooth, for simplicity we denote again $\mathcal{E}$ for the open subset $\mathcal{E}_{S}:=\tilde{f}_{X}^{-1}(S)\subset\mathcal{E}$,
  • $t\in S(\mathbb{C})$ is the point such that $\mathcal{Y}^{\prime}_{\mathbb{C},t}=Y_{0_{X}}$ and $\epsilon_{t}:E_{0}:=\mathcal{E}_{0,\mathbb{C},t}\to Y_{0_{X}}$ is surjective (there is at least one component of $E$ dominant over $Y_{O_{X}}$ since $E$ is dominant over $Y_{O_{X}}$ as $\mathcal{E}$ is dominant over $\mathcal{Y}^{\prime}_{\bar{S}}$), in particular $E_{0}$ is birational to $Y_{0_{X}}$,hence $E_{0}$ is birational to $X$. We denote $E_{i}:=\mathcal{E}_{i,\mathbb{C},t}$ for each $0\leq i\leq r$ and $E:=\mathcal{E}_{\mathbb{C},t}=\cup_{i=0}^{r}E_{i}$.
We then prove, assuming the Hodge conjecture for smooth projective varieties of dimension less or equal to $d-1$, the Hodge conjecture for the $E_{i}$, more specifically for $E_{0}$. Denote for $I\subset[0,\ldots,s]$, $i_{I}:\mathcal{E}_{I}\hookrightarrow\mathcal{E}$ and for $I\subset J\subset[0,\ldots,s]$, $i_{JI}:\mathcal{E}_{J}\to\mathcal{E}_{I}$ the closed embeddings. Then, $i_{\bullet}:\mathcal{E}_{\bullet}\to\mathcal{E}$ in $\mathop{\rm Fun}\nolimits(\Delta,\mathop{\rm Var}\nolimits(\mathbb{Q}))$ is a simplicial resolution with $\mathcal{E}_{\bullet}\in\mathop{\rm Fun}\nolimits(\Delta,\mathop{\rm SmVar}\nolimits(\mathbb{Q}))$. We have then
$E_{Hdg}(\mathcal{E}/S):=(\tilde{f}_{X}\circ i_{\bullet*}E_{zar}(\Omega^{\bullet}_{\mathcal{E}_{\bullet}/S},F_{b},W),(\tilde{f}_{X}\circ i_{\bullet*}E_{usu}(\mathbb{Q}_{\mathcal{E}_{\bullet,\mathbb{C}}^{an}},W)),\tilde{f}\circ i_{\bullet*}\alpha(\mathcal{E}_{\bullet}))\in D(MHM_{\mathbb{Q},gm}(S))$
Then, definition-proposition 1 applied to $f_{T}=f\times_{\mathbb{P}(l_{X})}\bar{T}^{o}:\mathcal{Y}^{\prime}_{\bar{T}^{o}}\to\bar{T}^{o}$ gives the distinguish in $D(MHM_{\mathbb{Q},gm}(S))$,
$E_{Hdg}(\mathcal{Y}^{\prime}_{S}/S)\xrightarrow{Sp_{\mathcal{Y}^{\prime}_{S}/f_{T}}:=(f_{T}\circ i_{\mathcal{Y}^{\prime}_{S}})_{*}Sp_{\mathcal{Y}_{S}/\mathcal{Y}^{\prime}_{\bar{T}^{o}}}}\psi_{S}E_{Hdg}(\mathcal{Y}^{\prime}_{T^{o}}/T^{o})$
$\xrightarrow{(f_{T}\circ i_{\mathcal{Y}^{\prime}_{S}})_{*}c(\psi_{\mathcal{Y}^{\prime}_{S}}\mathbb{Z}^{Hdg}_{\mathcal{Y}_{T^{o}}})}\phi_{S}E_{Hdg}(\mathcal{Y}^{\prime}_{T^{o}}/T^{o})\xrightarrow{(f_{T}\circ i_{\mathcal{Y}^{\prime}_{S}})_{*}c(\mathbb{Z}^{Hdg}_{\mathcal{Y}^{\prime}_{S})}}E_{Hdg}(\mathcal{Y}^{\prime}_{S}/S)[1]$
and definition-proposition 1 applied to $\tilde{f}_{X}:\mathcal{Y}_{\bar{T}^{o}}\to\bar{T}^{o}$ gives the distinguish in $D(MHM_{\mathbb{Q},gm}(S))$,
$E_{Hdg}(\mathcal{E}/S)\xrightarrow{Sp_{\mathcal{E}/\tilde{f}_{X}}:=(\tilde{f}_{X}\circ i_{\mathcal{E}})_{*}Sp_{\mathcal{E}/\mathcal{Y}}}\psi_{S}E_{Hdg}(\mathcal{Y}_{T^{o}}/T^{o})$
$\xrightarrow{(\tilde{f}_{X}\circ i_{\mathcal{E}})c(\psi_{\mathcal{E}}\mathbb{Z}^{Hdg}_{\mathcal{Y}_{T^{o}}})}\phi_{S}E_{Hdg}(\mathcal{Y}_{T^{o}}/T^{o})\xrightarrow{(\tilde{f}_{X}\circ i_{\mathcal{E}})_{*}c(\mathbb{Z}^{Hdg}_{\mathcal{E})}}E_{Hdg}(\mathcal{E}/S)[1]$
and $Sp_{\mathcal{E}/\tilde{f}_{X}}\circ\epsilon^{*}=Sp_{\mathcal{Y}_{\bar{T}^{o}}/f_{T}}$. Let $p\in\mathbb{Z}$. By the hard Lefchetz theorem on $H^{*}(E_{0})$, we may assume that $2p\leq d$. Taking the cohomology to this distinguish triangle, we get in particular, for each $p\in\mathbb{Z}$, the maps in $MHM_{\mathbb{Q},gm}(S)$
  • $Sp_{\mathcal{Y}^{\prime}_{S}/f_{T}}:E^{2p}_{Hdg}(\mathcal{Y}^{\prime}_{S}/S)\to H^{2p}\psi_{S}E_{Hdg}(\mathcal{Y}_{T^{o}}/T^{o})$
  • $Sp_{\mathcal{E}/\tilde{f}_{X}}:E^{2p}_{Hdg}(\mathcal{E}/S)\to H^{2p}\psi_{S}E_{Hdg}(\mathcal{Y}_{T^{o}}/T^{o})$.
Moreover, we have, see lemma 1, an isomorphism of variation of mixed Hodge structures over $S$
$c_{\mathcal{E}}:=H^{2p-1}(\tilde{f}_{X}\circ i_{\mathcal{E}})_{*}c(\mathbb{Q}^{Hdg}_{\mathcal{E}})\circ c_{V}:E_{Hdg,\mathcal{E}}^{2p-1}(\mathcal{Y}_{\bar{T}^{o}}/\bar{T}^{o})\xrightarrow{\sim}$
$\phi_{S}E_{Hdg,\mathcal{E}}^{2p-1}(\mathcal{Y}_{T^{o}}/T^{o})\xrightarrow{\sim}\ker(Sp_{\mathcal{E}/\tilde{f}_{X}}:E^{2p}_{Hdg}(\mathcal{E}/S)\to\psi_{S}E^{2p}_{Hdg}(\mathcal{Y}_{T^{o}}/T^{o})).$
For $\lambda\in H^{2p}(\mathcal{E}_{0,\mathbb{C},t^{\prime}}^{an},\mathbb{Q})$, $t^{\prime}\in S(\mathbb{C})$, we consider $\pi_{S}:\tilde{S}_{\mathbb{C}}^{an}\to S_{\mathbb{C}}^{an}$ the universal covering in $\mathop{\rm AnSm}\nolimits(\mathbb{C})$ and
$\mathbb{V}_{S}(\lambda):=\pi_{S}(\lambda\times\tilde{S}^{an}_{\mathbb{C}})\subset E^{2p,an}_{DR}(\mathcal{E}_{0,\mathbb{C}}/S_{\mathbb{C}})$
the flat leaf induced by the universal covering (see section 2.1) and
$\mathbb{V}_{S}^{p}(\lambda):=\mathbb{V}_{S}(\lambda)\cap F^{p}E^{2p}_{DR}(\mathcal{E}_{0,\mathbb{C}}/S_{\mathbb{C}})\subset E^{2p}_{DR}(\mathcal{E}_{0,\mathbb{C}}/S_{\mathbb{C}}),$
which is an algebraic variety by [6] (finite over $S$). Note that $\mathbb{V}^{p}_{S}(\lambda)\subset HL^{p,2p}(\mathcal{E}_{0,\mathbb{C}}/S_{\mathbb{C}})$ is the union of the irreducible components passing through $\lambda$. Similarly, for $\eta\in H^{2p}(\mathcal{E}_{\mathbb{C},t^{\prime}}^{an},\mathbb{Q})$, $t^{\prime}\in S(\mathbb{C})$, we consider $\pi_{S}:\tilde{S}_{\mathbb{C}}^{an}\to S_{\mathbb{C}}^{an}$ the universal covering in $\mathop{\rm AnSm}\nolimits(\mathbb{C})$ and
$\mathbb{V}_{S}(\eta):=\pi_{S}(\eta\times\tilde{S}^{an}_{\mathbb{C}})\subset E^{2p,an}_{DR}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})$
the flat leaf induced by the universal covering (see section 2.1) and
$\mathbb{V}_{S}^{p}(\eta):=\mathbb{V}_{S}(\eta)\cap F^{p}E^{2p}_{DR}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})\subset E^{2p}_{DR}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}}).$
We have
$H^{2p}E_{Hdg}(\mathcal{Y}_{T^{o}}/T^{o})=((E^{2p}_{DR}(\mathcal{Y}_{T^{o}}/T^{o}),F),R^{2p}\tilde{f}_{X*}\mathbb{Q}_{\mathcal{Y}^{an}_{T^{o},\mathbb{C}}},\tilde{f}_{X*}\alpha(\mathcal{Y}_{T^{o}}))\in MHM_{\mathbb{Q},gm}(T^{o})$
By [8], the filtered vector bundle
$(E^{2p}_{DR}(\mathcal{Y}_{T^{o}}/T^{o}),F)=H^{2p}\tilde{f}_{X*}E_{zar}(\Omega^{\bullet}_{\mathcal{Y}_{T^{o}}/T^{o}},F_{b})\in\mathop{\rm Vect}\nolimits_{\mathcal{D}fil}(T^{o})$
extend to a filtered vector bundle
$(E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o}}/T^{o}),F):=H^{2p}\tilde{f}_{X*}E_{zar}(\Omega^{\bullet}_{\mathcal{Y}_{T^{o}}/T^{o}}(\mathop{\rm log}\nolimits\mathcal{E}),F_{b})\in\mathop{\rm Vect}\nolimits_{\mathcal{D}fil}(\bar{T}^{o})$
such that
$H^{2p}\psi_{S}E_{Hdg}(\mathcal{Y}_{T^{o}}/T^{o})=(i_{S}^{*mod}(E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o}}/T^{o}),F,W),(\psi_{S}R^{2p}\tilde{f}_{X*}\mathbb{Q}_{\mathcal{Y}^{an}_{T^{o},\mathbb{C}}},W),\psi_{S}\tilde{f}_{X*}\alpha(\mathcal{Y}_{T^{o}}))$
and
$Sp_{\mathcal{E}/\tilde{f}_{X}}:=(Sp_{\mathcal{E}/\tilde{f}_{X}},Sp_{\mathcal{E}/\tilde{f}_{X}}):((E^{2p}_{DR}(\mathcal{E}/S),F,W),(R^{2p}(\tilde{f}_{X}\circ i_{\bullet})_{*}\mathbb{Q}_{\mathcal{E}^{an}_{\bullet,\mathbb{C}}},W),(\tilde{f}_{X}\circ i_{\bullet})_{*}\alpha(\mathcal{E}_{\bullet}))$
$\to(i_{S}^{*mod}(E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o}}/T^{o}),F,W)=((E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o}}/T^{o})\cap p^{-1}(S),F),W),$
$(\psi_{S}R^{2p}\tilde{f}_{X*}\mathbb{Q}_{\mathcal{Y}^{an}_{T^{o},\mathbb{C}}},W),\psi_{S}\tilde{f}_{X*}\alpha(\mathcal{Y}_{T^{o}}))$
where $p:E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o}}/T^{o})\to\bar{T}^{o}$ is the projection. For
$\alpha\in H^{2p}(\mathcal{Y}_{t^{\prime\prime}}^{an},\mathbb{Q})=i_{t^{\prime}}^{*}\psi_{S}R^{2p}\tilde{f}_{X*}\mathbb{Q}_{\mathcal{Y}_{T^{o},\mathbb{C}}^{an}},\,t^{\prime}\in S_{\mathbb{C}},\,t^{\prime\prime}\in T^{o}(\mathbb{C}),$
we consider
  • $\pi_{T^{o}}:\tilde{T}^{o,an}\to T_{\mathbb{C}}^{o,an}$ the universal covering in $\mathop{\rm AnSm}\nolimits(\mathbb{C})$, the factorization
    $\pi_{S}:\tilde{S}_{\mathbb{C}}^{an}\to\tilde{\bar{T}}_{\mathbb{C}}^{o,an}\xrightarrow{\pi_{\bar{T}^{o}}}\bar{T}_{\mathbb{C}}^{o,an}$
    and
    $\mathbb{V}_{T^{o}}(\alpha):=\pi_{T^{o}}(\alpha\times\tilde{T}_{\mathbb{C}}^{o,an})\subset E^{2p,an}_{DR}(\mathcal{Y}_{T^{o},\mathbb{C}}/T^{o}_{\mathbb{C}})$
    the flat leaf (see section 2.1) and
    $\mathbb{V}_{T^{o}}^{p}(\alpha):=\mathbb{V}_{T^{o}}(\alpha)\cap F^{p}E^{2p}_{DR}(\mathcal{Y}_{T^{o},\mathbb{C}}/T^{o}_{\mathbb{C}})\subset E^{2p}_{DR}(\mathcal{Y}_{T^{o}\mathbb{C}}/T^{o}_{\mathbb{C}}),$
    which is an algebraic subvariety (finite over $T^{o}$) by [6],
  • $\pi_{S}:\tilde{S}_{\mathbb{C}}^{an}\to S_{\mathbb{C}}^{an}$ the universal covering in $\mathop{\rm AnSm}\nolimits(\mathbb{C})$ and
    $\mathbb{V}_{S}(\alpha):=\pi_{S}(\alpha\times\tilde{S}^{an}_{\mathbb{C}})\subset H^{2p}\psi_{S}E_{DR}(\mathcal{Y}_{T^{o},\mathbb{C}}/T^{o}_{\mathbb{C}})^{an}=i_{S}^{*mod}(E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T_{\mathbb{C}}^{o}))^{an}$
    the flat leaf (see section 2.1) and
    $\mathbb{V}_{S}^{p}(\alpha):=\mathbb{V}_{S}(\alpha)\cap F^{p}H^{2p}\psi_{S}E_{DR}(\mathcal{Y}_{T^{o},\mathbb{C}}/T^{o}_{\mathbb{C}})^{an}\subset H^{2p}\psi_{S}E_{DR}(\mathcal{Y}_{T^{o}\mathbb{C}}^{/}T^{o}_{\mathbb{C}})^{an}.$
Consider the exact sequence of variation of mixed Hodge structure over $S$
$0\to W_{2p-1}E_{Hdg}^{2p}(\mathcal{E}/S)\to E_{Hdg}^{2p}(\mathcal{E}/S)\xrightarrow{(i_{i}^{*})_{0\leq i\leq r}:=q}\oplus_{i=0}^{r}E_{Hdg}^{2p}(\mathcal{E}_{i}/S)\to 0$
It induces the exact sequence of vector bundles over $S$
$0\to W_{2p-1}E_{DR}^{2p}(\mathcal{E}/S)\to E_{DR}^{2p}(\mathcal{E}/S)\xrightarrow{(i_{i}^{*})_{0\leq i\leq r}:=q}\oplus_{i=0}^{r}E_{DR}^{2p}(\mathcal{E}_{i}/S)\to 0,$
and the exact sequence of presheaves on $S_{\mathbb{C}}^{an}$,
[…]
where $HL^{2p,p}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})\subset E_{DR}^{2p}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})$, $HL^{2p,p}(\mathcal{E}_{i,\mathbb{C}}/S_{\mathbb{C}})\subset E_{DR}^{2p}(\mathcal{E}_{i,\mathbb{C}}/S_{\mathbb{C}})$ are the locus of Hodge classes. Let $\lambda\in F^{p}H^{2p}(E_{0}^{an},\mathbb{Q})$, where we recall $E_{0}=\mathcal{E}_{0,\mathbb{C},t}$ and $E=\mathcal{E}_{\mathbb{C},t}=\cup_{i=0}^{r}E_{i}$. Consider then $\tilde{\lambda}\in H^{2p}(E^{an},\mathbb{Q})$, such that $q(\tilde{\lambda})=\lambda$, and
$Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda})\in Sp_{E/f_{X}}(H^{2p}(E^{an},\mathbb{Q}))\subset i_{t}^{*}\psi_{S}R^{2p}\tilde{f}_{X*}\mathbb{Q}_{\mathcal{Y}_{T^{o},\mathbb{C}}^{an}}.$
Since the monodromy of $R^{2p}\tilde{f}_{X*}\mathbb{Q}_{\mathcal{Y}_{T^{o},\mathbb{C}}}^{an}$ is irreducible by Picard Lefchetz theory as $i_{T^{o}*}:\pi_{1}(T_{\mathbb{C}}^{o})\to\pi_{1}(T_{\mathbb{C}})\xrightarrow{\sim}\pi_{1}(\mathbb{P}(l_{X})\backslash\Delta)$ is surjective, we have
$Sp_{\mathcal{E}/\tilde{f}_{X}}(\ker q):=Sp_{\mathcal{E}/\tilde{f}_{X}}(W_{2p-1}E_{DR}^{2p}(\mathcal{E}/S))=0,$
using the fact that there exists a neighborhood $V_{\bar{S}\cup\bar{S}^{\prime}}\subset\bar{T}_{\mathbb{C}}$ of $\bar{S}_{\mathbb{C}}\cup\bar{S}^{\prime}_{\mathbb{C}}$ in $\bar{T}_{\mathbb{C}}$ for the usual complex topology such that the inclusion $i_{\bar{S}_{\mathbb{C}}\cup\bar{S}^{\prime}_{\mathbb{C}}}:\bar{S}_{\mathbb{C}}\cup\bar{S}^{\prime}_{\mathbb{C}}\hookrightarrow V_{\bar{S}\cup\bar{S}^{\prime}}$ admits a retraction $r:V_{\bar{S}\cup\bar{S}^{\prime}}\to\bar{S}_{\mathbb{C}}\cup\bar{S}^{\prime}_{\mathbb{C}}$ which is an homotopy equivalence. The locus of Hodge classes passing through $\lambda$
$\mathbb{V}_{S}^{p}(\lambda):=\mathbb{V}_{S}(\lambda)\cap F^{p}E^{2p}_{DR}(\mathcal{E}_{0,\mathbb{C}}/S_{\mathbb{C}})\subset E^{2p}_{DR}(\mathcal{E}_{0,\mathbb{C}}/S_{\mathbb{C}}),$
inside the De Rham vector bundle of $\tilde{f}_{X}\circ i_{E_{0}}:\mathcal{E}_{0}\to S$ thus satisfy
$\mathbb{V}_{S}^{p}(\lambda)=q(Sp_{\mathcal{E}/\tilde{f}_{X}}^{-1}(\mathbb{V}_{S}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))))\subset E^{2p}_{DR}(\mathcal{E}_{0\mathbb{C}}/S_{\mathbb{C}}),$ (2)
where
  • $\mathbb{V}_{S}(\lambda)\subset E^{2p}_{DR}(\mathcal{E}_{0,\mathbb{C}}/S_{\mathbb{C}})$, $\mathbb{V}_{S}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))\subset E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T_{\mathbb{C}}^{o})$, are the flat leaves, e.g. $\mathbb{V}_{S}(\lambda):=\pi_{S}(\lambda\times\tilde{S}^{an}_{\mathbb{C}})$ where $\pi_{S}:H^{2p}(E^{an},\mathbb{C})\times\tilde{S}_{\mathbb{C}}^{an}\to E^{2p,an}_{DR}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})$ is the morphism induced by the universal covering $\pi_{S}:\tilde{S}_{\mathbb{C}}^{an}\to S_{\mathbb{C}}^{an}$,
  • $q:=q\otimes\mathbb{C}:E^{2p}_{DR}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})\to E^{2p}_{DR}(\mathcal{E}_{0,\mathbb{C}}/S_{\mathbb{C}})$ is the quotient map.
  • $Sp_{\mathcal{E}/\tilde{f}_{X}}:=Sp_{\mathcal{E}/\tilde{f}_{X}}\otimes\mathbb{C}:E^{2p}_{DR}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})\to i_{S}^{*mod}E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T_{\mathbb{C}}^{o})$.
Now,
  • If $Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda})=0$, we have by lemma 1 applied to $\tilde{f}_{X}:\mathcal{Y}_{\bar{T}^{o}}\to\bar{T}^{o}$, $\tilde{\lambda}=c_{\mathcal{E}}(\tilde{\lambda})$ with $\tilde{\lambda}\in F^{p}H^{2p-1}_{E}(\mathcal{Y}^{an}_{\mathbb{C},C},\mathbb{Q})$ where $C\subset\bar{T}_{\mathbb{C}}^{o}$ is a smooth transversal slice, hence $\mathop{\rm dim}\nolimits(C)=1$ and $\mathcal{Y}_{C}$ is smooth, such that $C\cap S_{\mathbb{C}}=\left\{t\right\}$ and
    $\mathbb{V}^{p}_{S}(\tilde{\lambda})=c_{\mathcal{E}}(\mathbb{V}^{p}_{S}(\tilde{\lambda})),\;\mathbb{V}^{p}_{S}(\tilde{\lambda})\subset E_{DR,\mathcal{E}_{\mathbb{C}}}^{2p-1}(\mathcal{Y}_{\bar{T}^{o},\mathbb{C}}/\bar{T}_{\mathbb{C}}^{o})$
    with
    $c_{\mathcal{E}}=(c_{\mathcal{E}},c_{\mathcal{E}}):((E_{DR,\mathcal{E}}^{2p-1}(\mathcal{Y}_{\bar{T}^{o}}/\bar{T}^{o}),F,W),(R^{2p-1}\tilde{f}_{X*}\circ\Gamma_{\mathcal{E}}\mathbb{Q}_{\mathcal{Y}_{\bar{T}^{o},\mathbb{C}}^{an}},W),\alpha(\mathcal{Y}_{\bar{T}^{o}}))$
    $\to((E_{DR}^{2p}(\mathcal{E}/S),F,W),(R^{2p}(\tilde{f}_{X}\circ i_{\mathcal{E}})_{*}\mathbb{Q}_{\mathcal{E}_{\mathbb{C}}^{an}},W),\alpha(\mathcal{E}_{\bullet}))$
    We have for each $t^{\prime}\in S(\mathbb{C})$, the isomorphism
    $\oplus_{cardI=2}F^{p}H^{2p}(E^{an}_{I\mathbb{C},t^{\prime}},\mathbb{Q})=F^{p}\mathop{\rm Gr}\nolimits^{W}_{2p}H^{2p-2}((\bar{\mathcal{Y}_{\mathbb{C},C^{\prime}}}\backslash E)^{an},\mathbb{Q})\xrightarrow{c(\mathbb{Z}(\mathcal{Y}_{C}\backslash\mathcal{E}_{\mathbb{C},t^{\prime}}))}$
    $F^{p}\mathop{\rm Gr}\nolimits^{W}_{2p}H^{2p-1}_{E}(\bar{\mathcal{Y}^{an}_{\mathbb{C},C^{\prime}}},\mathbb{Q})=F^{p}\mathop{\rm Gr}\nolimits^{W}_{2p}H^{2p-1}_{E}(\mathcal{Y}^{an}_{\mathbb{C},C^{\prime}},\mathbb{Q})$
    where $C\subset\bar{T}_{\mathbb{C}}^{o}$ is a smooth transversal slice, hence $\mathop{\rm dim}\nolimits(C^{\prime})=1$ and $\mathcal{Y}_{C^{\prime}}$ is smooth, such that $C^{\prime}\cap S_{\mathbb{C}}=\left\{t^{\prime}\right\}$ and $\bar{\mathcal{Y}_{C^{\prime}}}\in\mathop{\rm PSmVar}\nolimits(\mathbb{C})$ is a smooth compactification of $\mathcal{Y}_{C^{\prime}}$, the last equality follows from excision (note that $E$ is proper). Hence, assuming the Hodge conjecture for smooth projective varieties of dimension less or equal to $d-1$,
    $\mathbb{V}^{p}_{S}(\lambda)\subset\mathop{\rm Gr}\nolimits^{W}_{2p}E_{DR,\mathcal{E}_{\mathbb{C}}}^{2p-1}(\mathcal{Y}_{\bar{T}^{o},\mathbb{C}}/\bar{T}_{\mathbb{C}}^{o}),\;\mathbb{V}^{p}_{S}(\lambda)=c_{\mathcal{E}}(\mathbb{V}^{p}_{S}(\lambda))\subset E^{2p}_{DR}(\mathcal{E}_{0\mathbb{C}}/S_{\mathbb{C}})\subset\mathop{\rm Gr}\nolimits^{W}_{2p}E^{2p}_{DR}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})$
    are defined over $\bar{\mathbb{Q}}$ and its Galois conjugates are also components of the locus of Hodge classes, since $\mathop{\rm dim}\nolimits(E_{I})=d-1$ for $I\subset[0,\cdots,r]$ such that $cardI=2$.
  • Consider now the case where $Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda})\neq 0$. We have the key equality
    $\mathbb{V}_{S}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))=\overline{\mathbb{V}_{T^{o}}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))}\cap p^{-1}(S_{\mathbb{C}})\subset E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T^{o}_{\mathbb{C}})$
    where $p:=p\otimes\mathbb{C}:E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T^{o}_{\mathbb{C}})\to\bar{T}^{o}_{\mathbb{C}}$ is the projection, the inclusion $\subset$ is obvious whereas the inclusion $\supset$ follows from [6] lemma 2.11 which state the invariance of $\tilde{\lambda}$ under the monodromy
    $\mathop{\rm Im}\nolimits(\pi_{1}(p(\mathbb{V}_{T^{o}}^{p}(\tilde{\lambda}))\cap D^{*})\to\pi_{1}(D^{*})),\;(D,D^{*})\to(\bar{T}^{o,an}_{\mathbb{C}},\bar{T}^{o,an}_{\mathbb{C}}\backslash S_{\mathbb{C}}^{an})).$
    But by [5] theorem 1, the Hodge conjecture holds for projective hypersurfaces hence
    $\mathbb{V}_{T^{o}}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))\subset E^{2p}_{DR}(\mathcal{Y}_{T^{o},\mathbb{C}}/T^{o}_{\mathbb{C}})$
    is an algebraic subvariety defined over $\bar{\mathbb{Q}}$ and its Galois conjugates are also components of the locus of Hodge classes. Hence,
    $\mathbb{V}_{S}^{p}(\lambda)=q(Sp_{\mathcal{E}/\tilde{f}_{X}}^{-1}(\mathbb{V}_{S}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))))\subset E^{2p}_{DR}(\mathcal{E}_{0\mathbb{C}}/S_{\mathbb{C}})$
    is an algebraic subvariety defined over $\bar{\mathbb{Q}}$ and its Galois conjugates are also components of the locus of Hodge classes.
The fact that $\mathbb{V}_{S}^{p}(\lambda)\subset E^{2p}_{DR}(\mathcal{E}_{0\mathbb{C}}/S_{\mathbb{C}})$ is defined over $\bar{\mathbb{Q}}$ and its Galois conjugates are also components of the locus of Hodge classes gives, by [4] theorem 4,
$\lambda=[Z]\in H^{2p}(E^{an}_{0},\mathbb{Q}),\,\mbox{with}\,Z\in\mathcal{Z}^{p}(E_{0}).$
Since $p\in\mathbb{Z}$ and $\lambda\in F^{p}H^{2p}(E_{0}^{an},\mathbb{Q})$ are arbitrary, the Hodge conjecture holds for $E_{0}$. Proposition 1(ii) and the fact that the Hodge conjecture is true for smooth complex projective varieties of dimension less or equal to $d-1$ by induction hypothesis then implies that Hodge conjecture holds for $X$, since $X$ is birational to $E_{0}$ and $X,E_{0}\in\mathop{\rm PSmVar}\nolimits(\mathbb{C})$ are connected of dimension $d$. ∎